Opened 5 years ago

Closed 3 years ago

#536 closed flight processing (fixed)

GB13/01, flight day 221/2014, Bury St Edmunds

Reported by: knpa Owned by: benj
Priority: alpha 4 high Milestone: 2014 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/GB13_01-2014_221_Bury_St_Edmunds/

Data arrived from ARSF via network transfer on 14/08/2014

GB13/01 Scientific objective: a combination of an ecological model with remotely sensed data can result in an accurate, high resolution map of vegetation communities. Such maps can form the basis to model the distribution of animal species or groups.

GB13/01 Priority: 9

PI: Tim Benton

Notes: There is FW A couple of lines are very long Nav is in two parts, it restarts after first line

Sensors:

Camera (Requested) Delivered 11/11/2014
Fenix (Requested) Delivered 11/11/2014
LiDAR (Requested) Delivered 11/11/2014
FW LiDAR (Not Requested) Delivered 11/11/2014

Change History (37)

comment:1 Changed 5 years ago by tec

RCD
Started converting Raw images to TIFF

Last edited 5 years ago by tec (previous) (diff)

comment:2 Changed 5 years ago by tec

RCD
Removing image 09083820100001G4.jpg

Waiting for Nav to continue.

Last edited 5 years ago by tec (previous) (diff)

comment:3 Changed 5 years ago by tec

NAV
Started nav processing.

comment:4 Changed 5 years ago by tec

ATTL Basestation Data
Latitude 52 31 00.58336
Longitude 0 59 17.28458
El Height81.486m


NEWR Basestation Data
Latitude 52 16 00.50332
Longitude 0 24 54.96024
El Height75.831m
Last edited 5 years ago by tec (previous) (diff)

comment:5 Changed 5 years ago by tec

Nav
Nav done, copying back now, elevation mask of 16 used.

comment:6 Changed 5 years ago by tec

Nav
Finished Nav

comment:7 Changed 5 years ago by tec

RCD
Finished RCD processing. Awaiting DC

comment:8 Changed 5 years ago by lah

Delivery Check Complete
All fine, except missing first 14 entries from eventfile. These have purposely been removed by tec and mark1, so ready to deliver.

comment:9 Changed 5 years ago by dap

Beginning Hyperspectral Fenix Processing

comment:10 Changed 5 years ago by dap

Fenix Processing

Moved all Fenix files for line 1 to <proj_directory>/hyperspectral/fenix/line_1 as the nav crashed and the navigation doesn't cover it to generate a dem.

comment:11 Changed 5 years ago by dap

Fenix Processing

Copied all raw header files to hyperspectral/fenix/<original file name>.orig and edited original raw header files so wavelengths and fwhm are same as those in the raw calibration header.

comment:12 Changed 5 years ago by dap

Fenix Processing

Found SCT values as follows:

Flightline FENIX
2 0.90
3 0.99
4 0.94
5 0.95
6 0.97

comment:13 Changed 5 years ago by tec

LiDAR Processing
Starting LiDAR processing

comment:14 Changed 5 years ago by dap

Fenix Processing

Created delivery, now ready for delivery check.

comment:15 Changed 5 years ago by tec

LiDAR Processing
Line 115 at 09:01 will not process in alspp but was reflown at 10:20

Roll Errors

Flightline Roll Error
090108 N/A
092723 -0.006375
093713 -0.006375
095026 -0.006300
100431 -0.006300
102044 -0.006300

comment:16 Changed 5 years ago by tec

LiDAR Processing
Creating a delivery.

comment:17 Changed 5 years ago by tec

LiDAR Processing
Finished creating delivery, ready for delivery check.

comment:18 Changed 5 years ago by lah

Starting Fenix Delivery Check

comment:19 Changed 5 years ago by lah

Fenix DC complete
Some changes to the ReadMe required:

  • PI doesn't agree with logsheet, but does with ticket, which is correct?
  • Pg9 dem name needs correcting _ instead of -
  • Underflows need adding

All lines suffer bad underflows between 410-444 and 488-620
Line 1 has underflows at band 348 and 374-380
Line 2 320-350
Line 3 has some underflows 444-488
Line 5 345-353, 373-382

I have changed the screenshots names as were from 2-6, added crlf terminator to txt file, removed extra 221 from proj_inf xml and changed hyp_genreadme link - to _ to run check_apl_cmd (still needs changing in readme). Evrything else is fine.

comment:20 Changed 5 years ago by lah

Starting Lidar Delivery Check

comment:21 Changed 5 years ago by dap

Fenix Delivery Check

  • Changed PI in logsheet to 'Tim Benton' after checking confirmation email from ops when data was unpacked for correct PI.
  • Changed dem file name in example aplcorr command on page 9 of read me to match the name of the dem in the dem/ directory.
  • Added table in data quality remarks section of read me to give information about which bands contain underflows.

Fenix now ready for delivery or rechecking.

comment:22 Changed 5 years ago by lah

Fenix
Copied across ReadMe pdf with formatted table and zipped mapped flight lines. Ready for delivery.

comment:23 Changed 5 years ago by lah

Lidar DC complete
Some issues to check:

  • Many missing waveforms and causes seek error with waveviewer (except line 5), though all zipped fine.
  • Some missed noise, particularly parts of powerlines. Probably ok, but worth taking a little extra time in future as these don't take long to make consistent.
  • Read Me is ok, but might be work making consistent with Fenix and using 115b instead of just 115. Also should be overlap differences.

The dem is definitely too big, but otherwise fine. demcompare returned means of -1.28 and -0.023 (unmasked). I've changed header to 2.0.

I have copied over the logsheet with the correct PI and crlf'd the txt files.

comment:24 Changed 5 years ago by tec

Waveforms should be fine.
Steve said dont worry about powerlines but will be more careful next time.
Changed 115 to 115b

Dem should be fine, will wait till mark1 gets back.

comment:25 Changed 5 years ago by tec

Mark1 says dem is fine.

Shall mark ready to be delivered as its only minor changes.

comment:26 Changed 5 years ago by tec

Delivery
RCD, Fenix and LiDAR delivered to Tim on 11/11/2014

comment:27 Changed 5 years ago by tec

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:28 Changed 4 years ago by mark1

Fenix
Fenix data needs reprocessing due to fault with sensor. Incorrect SWIR times recorded in raw files. Have changed tint2 from 13.7 to 2.15 in all raw hdr files.

comment:29 Changed 4 years ago by mark1

Fenix data have been reprocessed.

comment:30 Changed 4 years ago by lah

Fenix re-DC
VNIR now looks too high compared to SWIR. mark1 looking into.

comment:31 Changed 4 years ago by dap

Fenix Delivery Check Take 3

Finished delivery check. Vegetation spectra look similar to examples on Delivery Checking procedure page. Only other issue was the broken link to aster policies.

comment:32 Changed 4 years ago by dap

Fenix

Have recreated readme with ASTER citations and fixed link to 'citing our data' page.
It appears that someone else has zipped the mapped files, so new delivery on molucca2 is ready for delivery.

Last edited 4 years ago by dap (previous) (diff)

comment:33 Changed 4 years ago by dac

Fenix

Reprocessed data delivered to Tim Benton on USB harddrive 31/07/2015

comment:34 Changed 4 years ago by dac

Fenix Delivery

Hard drive received back

comment:35 Changed 4 years ago by asm

Archiving - Still pending.

Last edited 4 years ago by asm (previous) (diff)

comment:36 Changed 3 years ago by dac

Archiving

Started uploading data to NEODC.

comment:37 Changed 3 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.